OpenStudy (anonymous):

If g(x)=(x-1)^2, what is g(t+at)

OpenStudy (anonymous):

\[g(t+\Delta t)\]

OpenStudy (anonymous):

g(t+at)=g(t+at-1)^2

OpenStudy (anonymous):

\[g(x)=x^2-2x+1\] \[g(x+\Delta x)=(x+\Delta x)^2+(x+\Delta x)+1\] then some algebra

OpenStudy (anonymous):

\[g(x+Delta x)=x^2+2x\Delta x+(\Delta x)^2+x+\Delta x +1\]

OpenStudy (anonymous):

g(t+at)=t^2+2at^2+a^2t^2-2at-2t